2- A sample of 49 sudden infant death syndrome (SIDS) cases had a mean birth weight of 2998 g. Based on other births in the county, we will assume ơ 800g. Calculate the 95% confidence interval for the mean birth weight of SIDS cases in the county. Interpret your results.
Answer #1

Solution:

Confidence Interval for population mean using z distribution

Given,

\bar x = 2998 ....... Sample mean

\sigma   = 800 ........Population standard deviation

n = 49 ....... Sample size

Note that, Population standard deviation(\sigma) is known..So we use z distribution.

Our aim is to construct 95% confidence interval.

\therefore c = 0.95

\therefore\alpha = 1- c = 1- 0.95 = 0.05

\therefore  \alpha/2 = 0.05 \slash 2 = 0.025 and 1- \alpha /2 = 0.975

Search the probability 0.975 in the Z table and see corresponding z value

\thereforeZ_{\alpha/2} = 1.96 for the 95% Confidence interval

Now , confidence interval for mean(\mu) is given by:

\bar x - Margin. of .error < \mu < \bar x + Margin. of .error

  \bar x-Z_{\alpha/2}*(\sigma/\sqrt{n}) < \mu < \bar x+Z_{\alpha/2}*(\sigma/\sqrt{n})

2998 -1.96*(800/ \sqrt{} 49)   < \mu <    2998 +1.96*(800/ \sqrt{} 49)

2998-224 < \mu <   2998+224

2774 < \mu < 3222

So, (2774 , 3222) is the required 95% confidence interval for mean.

95% C.I. means that there is 0.95 probability that the true mean (population mean) will lie within the interval.